OpenStudy (anonymous):

Place the indicated product in the proper location on the grid. (4ya - 6xb)(4ya + 6xb)

OpenStudy (confluxepic):

Ok.

OpenStudy (anonymous):

I don't really understand your question What grid? (4ya - 6xb)(4ya + 6xb} = 16(ya)² -36(xb)² \[\frac{ y^2 }{ (1/4)^2 }-\frac{ x^2 }{ (1/6)^2 }\] if you would put it on a grid with ya axis as an y-axis and xb as an x-axis and you equate it to a value, say 1 you get an hyperbola See attachment for \[\frac{ y^2 }{ (1/4)^2 }-\frac{ x^2 }{ (1/6)² }=1\]